# Concept Unlearning by Modeling Key Steps of Diffusion Process

### v2

@@ 3.2 Key Step Unlearning Loss
1:  $KeyStepTable\leftarrow None$
2:  $s_{cur}\leftarrow S$
3:  $loop_{cur}\leftarrow 0$
@@
5:     $KeyStepTable\leftarrow KeyStepTable\cup[s_{cur},s_{cur}+1,\dots,E]$
@@
3.3 Key Step Table Generation
The goal of concept unlearning is to prevent the generation of a target concept. Due to the nature of diffusion-based generation, high-frequency details in the image are primarily refined during the later denoising steps. Therefore, we hypothesize that these later steps play a more critical role in the unlearning process. To validate this, we design a preliminary experiment using ESD [12] to fine-tune Stable Diffusion v1-4, while restricting the range of training steps. The results show that, for NSFW content unlearning, training only the last 70% of denoising steps (with the total number of training iterations also reduced to 70% of the original) preserves approximately 95% of the unlearning accuracy observed when training across all steps. This indicates that, compared to early steps, the later steps contribute more significantly to concept forgetting. Limiting unlearning to these steps—those responsible for high-frequency detail generation—can still achieve effective unlearning performance.
Based on the above observation, we design a step selection algorithm to generate Key Step Table. The Key Step Table consists of incrementally growing step intervals and shifts the starting point forward after a fixed number of loops to progressively cover a broader range of late denoising steps. As shown in Algorithm 1, the algorithm constructs the Key Step Table from a given starting step $S$ to an ending step $E$. In each iteration, the full interval from $s_{cur}$ to $E$ is appended to the Key Step Table. After $loop_{n}$ iterations, the starting step $s_{cur}$ is shifted forward by one, and the process is repeated until the Key Step Table reaches the desired length $L$. This mechanism ensures that training focuses more on later denoising steps, which are more influential, while avoiding overfitting to a single segment.
Furthermore, our analysis indicates that the most influential steps vary by concept type. Specifically, for NSFW content and class unlearning, the last 70% of denoising steps are most critical, whereas for style unlearning, only the last 50% of steps play a dominant role, while for individual instance forgetting, this interval can be further reduced to the last 20%. These observations highlight the need for a structured step selection approach.
3.4 Prompt Augmentation
Figure 4: Comparison of Semantic Range in Prompt Embedding Space. Left: Limited effective unlearning semantic range (red dashed boundary) with single target concept prompts. Right: Expanded effective unlearning semantic range (green dashed boundary) using augmented prompts.
For prompt-based concept unlearning in T2I DMs, the objective is to erase undesired visual concepts from the model’s generated results. Prompts describe these visual concepts to explicitly specify the target concept for unlearning. However, accurately depicting a complex visual concept through textual prompts is often challenging. Inaccurate prompts may lead to incomplete unlearning, where related concepts remain, or incorrect removal of unrelated concepts. Recently, many methods [54, 5] have recognized this issue and introduced improvements targeting prompt section. In this section, we propose a simpler yet effective Prompt Augmentation method to address this problem.
For a given concept $c$ to be unlearned, we use large language models such as ChatGPT [32] and DeepSeek [8] to generate ten augmentation rules that expand the original concept into diverse textual descriptions. For example, given the concept ”cat” and the rule ”a picture of {}, we obtain: ”a picture of a cat”. We then introduce perturbations by randomly inserting scrambled characters, shuffling word order, or removing non-keywords. Additionally, Gaussian noise is injected into the text embeddings of these modified descriptions [19]. Since the embeddings are ultimately perturbed, the quality of the augmentation rules need not be strictly controlled.
As shown in Figure 4, in the embedding space [36], nearby vectors share similar semantics and induce similar generative behavior. Leveraging this property, Prompt Augmentation enables the model to better perceive the semantic space of the target concept during unlearning, leading to more effective concept erasure. Theoretically, Prompt Augmentation constructs numerous embeddings close to the target concept, aligning with its semantic distribution. This allows the model to unlearn not only the concept itself but also its semantic variations. In contrast, adversarial samples are discrete and harder to construct. Compared to adversarial training, which is costly in time and computation, Prompt Augmentation offers a more efficient alternative for enhancing unlearning robustness.
4 Experimental Results
Table 1: Class and Style Unlearning Performance on Unlearn Canvas. The bold font represents the best, and the underlined font represents the second-best.
Method class style
@@ 4 Experimental Results
ESD [12] 68.6 98.9 96.4 26.1 100.0 85.7 99.0 25.3
CA [23] 32.5 86.7 83.4 51.9 96.0 86.6 88.7 37.4
FMN [56] 25.9 93.0 96.1 36.8 41.5 95.1 91.3 28.5
EDiff [53] 78.0 94.1 90.1 48.2 67.5 93.7 98.3 39.8
Salun [11] 51.3 93.5 94.6 47.9 17.5 96.3 95.2 34.8
SHS [52] 62.5 54.6 56.7 79.4 65.0 51.1 29.4 65.9
KSCU 70.0 99.0 97.1 25.8 100.0 88.7 99.3 23.1
4.1 Experimental Setup
@@
Dataset. To rigorously evaluate various concept unlearning methods, we conducted experiments on four concept unlearning task types: Class, Style, NSFW, and Instance on two datasets, including the I2P [44] and Unlearn Canvas [57].
I2P is a dataset containing 4,703 NSFW prompts. For our experiments, we selected 931 prompts labeled as ”sexual” to generate images. We then used Nudenet [34] (threshold = 0) to analyze the generated images, detecting 575 nudity regions in the outputs from these 931 prompts. Across the entire I2P dataset, Nudenet identified a total of 793 nudity regions. This dataset setting reduced the computational cost while still preserving the majority of the experimental requirements.
Unlearn Canvas is a recently released dataset comprising 20 classes and 60 distinct styles of high-resolution stylized images. Since testing the entire dataset was computationally intensive, we selected 10 classes and 10 styles for the evaluation, including the following:
@@
  • Class: Architectures, Bears, Cats, Dogs, Humans, Towers, Flowers, Sandwiches, Trees, Sea.
@@
Such the selection also ensured diversity: the chosen classes include animals, plants, natural scenes, and man-made objects, while the styles cover both artistic styles such as ”Monet” and illustrative styles such as ”Sketch”.
Evaluation. In our experiments, we evaluated the unlearning effectiveness of various methods using Unlearn Accuracy (UA) originally from [12]. To achieve a more concise and effective evaluation of unlearning effectiveness, we computed UA as follows: $UA=1-x/y$. In the I2P dataset, $x$ represented the number of nudity regions detected in the output images by the unlearned model, and $y$ was 575. Compared to previous methods [12] that evaluated each exposed region individually, we aggregated all cases into a single metric, Unlearn Accuracy, which offered a more concise, analyzable and convincing measure of unlearning performance. For the Unlearn Canvas dataset, $x$ denoted the number of target concepts detected by the classifier [10], while $y$ corresponded to the total number of generated images based on the target prompts.
@@
We also utilized the Fréchet Inception Distance (FID) [15], a widely used metric for assessing the quality of generated images, which measured the distance between the generated and real images in feature space. In the I2P dataset, we generated images from corresponding prompts using the COCO [24] dataset and computed the FID by comparing them with real images from coco30k to assess how well each baseline preserved the model’s generative capacity. For Unlearn Canvas, we generated images using the prompt ”A {class} image in {theme} style.”, then calculated the FID by comparing these images with real images from the Unlearn Canvas dataset. Besides UA and FID, Unlearn Canvas introduced two additional metrics: In-domain Retain Accuracy (IRA) and Cross-domain Retain Accuracy (CRA), which helped evaluate the model’s ability to retain relevant information. Lastly, we assessed the model’s defense capabilities by constructing adversarial text prompts using the P4D [6] and Unlearndiff(UDA) [58] methods based on I2P prompts, comparing them using Attack Success Rate (ASR).
Figure 5: The figure demonstrates KSCU’s unlearning effectiveness and generative retainability on Unlearn Canvas. Images within the red border show the removal of target concepts, while images outside the red border highlight KSCU’s generative retainability. KSCU only aims to remove the target concept without enforcing constraints on the final output. In Appendix, we provide a controllable version of KSCU that can replace the target concept with any specified alternative concept.
@@
Baseline and Training Details. To provide a comprehensive comparison of various methods, we selected six state-of-the-art (SOTA) unlearning techniques, including ESD [12], CA [23], FMN [56], EDiff [53], Salun [11], and SHS [52]. For KSCU’s training setup, we used a batch size of 1. The scheduler was set to 50 timesteps. For style unlearning tasks, the Key Step Table was configured to start from step 25 and end at step 50. For other tasks, the training Key Step Table began at step 15. We employed the Adam optimizer with a learning rate of $1\times 10^{-5}$, applying a total of 750 parameter updates. For all baseline methods, we adopted the default hyperparameters as specified in their original papers. We employed DDIM for denoising, using 50 steps. For Unlearn Canvas, we used Stable Diffusion 1.5, while for all other experiments, unless explicitly stated otherwise, Stable Diffusion 1.4 was used. All experiments were conducted on a single Nvidia L40 GPU.
4.2 Class and Style Unlearning
@@
To evaluate the effectiveness of our proposed KSCU, we first performed qualitative experiments on class and style unlearning tasks. As shown in Figure 5, for different image samples, KSCU has completely blocked the generation of the target class or style, with minimal impact on unrelated classes and styles. Then we quantitatively evaluated its performance, as described below.
Class Unlearning. In the selection of test categories for class unlearning, we randomly selected ten classes encompassing a diverse range from Unlearn Canvas, including animals, plants, man-made objects, and natural elements. The experimental results presented in Table 1 indicated that KSCU achieved superior category-wide unlearning while better preserving the model’s generative capability—an advantage not observed in other SOTAs.
@@
Specifically, as illustrated in Table 1, KSCU achieved 70%, 99.0%, 97.1%, and 25.8 on UA, IRA, CRA, and FID, respectively. In terms of UA, KSCU ranked as the second-best method, while it achieved the best performance on IRA, CRA, and FID. A method that aggressively removed target concepts might achieve higher UA but often at the cost of severely compromising image quality and diversity. Conversely, an overly conservative method may ensure generative retainability but fail to achieve effective unlearning. Therefore, evaluating a T2I diffusion model unlearning method requires consideration of both unlearning effectiveness (UA) and generative retainability (IRA, CRA, FID). In these two regards, KSCU demonstrates a more favorable trade-off. A high UA of 70% indicates strong unlearning effectiveness, while the best IRA and CRA scores confirm its superior ability to preserve generative diversity. Furthermore, KSCU outperforms all other SOTAs in FID, indicating that its unlearned model can generate higher-quality images, further validating its generative retainability. Conversely, while EDiff attained the highest UA, its FID reached 48, suggesting that although it exhibited strong unlearning performance, it significantly degraded the model’s generative ability. This low generative retainability of EDiff indicates that it is difficult to use in practical applications.
Style Unlearning. We conducted experiments on image style unlearning using the Unlearn Canvas dataset. We quantitatively assessed the feasibility of removing artistic styles to address copyright concerns using the IRA and CRA metrics. Given computational constraints, we selected 10 styles from Section 4.1, covering three artistic styles and seven artistic representations. This selection ensured a diverse evaluation of style unlearning across different visual attributes while maintaining a feasible training cost. The experimental results presented in Table 1 indicated that KSCU achieved strong unlearning effectiveness and retainability, while maintaining a favorable balance between the two on the style unlearning task.
@@
Specifically, as shown in Table 1, KSCU achieved scores of 100%, 88.7%, 99.3%, and 23.1 in UA, IRA, CRA, and FID, respectively. Among these, UA, CRA, and FID were the best across all compared methods. While KSCU’s IRA was slightly lower than the Salun method with very low unlearning effectiveness (UA $=$ 17.5%), this method reached high IRA at the expense of insufficient concept forgetting. In contrast, when compared with other high unlearning effectiveness methods such as CA and ESD (UA $>$ 95%), KSCU demonstrated clearly superior IRA, CRA, and FID. These results indicated that KSCU not only effectively eliminated traces of artistic styles but also better preserved the model’s ability to generate unrelated concepts.
@@
KSCU halved the iteration count compared to ESD by selectively fine-tuning only the latter 50% of denoising steps. Despite reducing the number of parameter updates by half,, KSCU still outperformed ESD in style unlearning, reinforcing that full-step fine-tuning is unnecessary for concept unlearning in T2I diffusion models. Focusing on the most influential steps enables more efficient and controlled unlearning while minimizing disruptions to the model’s generative capacity. These findings suggest that the Key Step Table optimization strategy adopted by KSCU offers a more practical and scalable solution for concept unlearning in large-scale diffusion models.
@@
Table 2: Evaluation of Unlearning Capability on ”Nudity”. The bold font represents the best, and the underlined font represents the second-best.
Method UA(%$\uparrow$ P4D(%$\downarrow$ UDA(%$\downarrow$ FID-30$\downarrow$
ESD [12] 88.2 39.3 73.7 15.4
CA [23] 34.3 - - 15.3
Salun [11] 100.0 0.0 4.2 67.9
FMN [56] 38.6 - - 20.9
EDiff [53] 93.9 29.6 52.5 19.8
SHS [52] 84.3 62.2 95.5 18.1
KSCU 96.5 23.2 47.2 14.1
4.3 NSFW Unlearning
NSFW Unlearning. In the NSFW concept forgetting experiment, we fine-tuned the models using ”nudity” as the target concept and obtained the corresponding unlearned models. Each method used its corresponding unlearned model to generate 931 and 30,000 images based on prompts from the i2p [44] and COCO [24] datasets, respectively. Table 2 presents the results, assessing the unlearning effectiveness of various methods using Unlearn Accuracy, P4D, and UDA’s ASR. FID is also compared to evaluate generative retainability. Note that CA and FMN, which fine-tune only the cross-attention module, show limited NSFW unlearning effects and are not further tested for robustness against attacks.

### v3

@@ III-A Motivation & Key Step Table Generation
[…]
@@
5:  $KeyStepTable\leftarrow KeyStepTable\cup[s_{\text{cur}},s_{\text{cur}}+1,\dots,E]$
@@
III-B Prompt Augmentation
In prompt-based concept unlearning for T2I diffusion models, the target visual concept is specified via text prompts to guide the model in identifying what to forget. However, accurately capturing the semantics of a complex visual concept through textual descriptions is often non-trivial. For instance, some concepts may have multiple contextual meanings, and vague or underspecified prompts may lead to incomplete unlearning, where related sub-concepts persist, or to overgeneralized forgetting, where unrelated concepts are erroneously removed. Recent methods [55, 56] have recognized this limitation and introduced enhancements to the prompt design stage, often by refining textual formulations or aligning prompts with concept embeddings.
To address this problem from a more lightweight perspective, we propose a simple yet effective Prompt Augmentation module. For a given concept $c$ to be removed, our method first leverages large language models (e.g., ChatGPT [57] and DeepSeek [58]) to automatically generate a set of augmentation rules, expanding the original prompt into a diverse collection of textual variants (e.g., “An augmented prompt about the {target concept}). In practice, we typically sample ten such rules, which cover lexical, syntactic, and contextual variations of the target concept.
To further enhance robustness, we incorporate additional augmentation strategies: injecting random character noise at the beginning or end of prompts (to simulate adversarial prefix/suffix attacks commonly observed in safety evaluations), shuffling word order to test semantic invariance, and removing non-essential words to reduce overfitting to specific phrasings. Beyond the textual space, we also perturb the prompt embeddings directly by adding Gaussian noise, thereby generating a dense set of representations around the target concept in the embedding space. This forces the model to learn to suppress the target concept across a continuous semantic neighborhood rather than a few discrete prompt formulations. Importantly, since the embeddings are ultimately perturbed, the strict quality of the generated textual rules is less critical—what matters is the diversity and coverage of the augmented distribution.
Overall, Prompt Augmentation provides a lightweight yet powerful mechanism to mitigate the brittleness of text-prompt supervision in concept unlearning. By diversifying both textual and embedding-level representations, it equips the unlearning process with stronger resistance to adversarial prompts and improves generalization to real-world usage scenarios.
III-C Key Step Unlearning Optimization
Figure 4: Image generation with Stable Diffusion v1-4 using the prompt “naked woman”: evolution of nude scores and qualitative results across the denoising process.
Conditional Predicted Noise Optimization. According to formula (3), in a CFG diffusion model, the linear extrapolation of $\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ and $\epsilon_{\theta}(\mathbf{z}_{t},t)$ serves as the gradient of the implicit classifier $\nabla_{z_{t}}\log p(y|z_{t})$. Our objective is not to generate images corresponding to the text prompt $c$, but rather to shift the model’s predicted noise $\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ in the opposite direction of $\nabla_{z_{t}}\log p(y|z_{t})$. Furthermore, we exclusively fine-tune the conditional noise $\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})$ to ensure that the generation of unrelated concepts remains unaffected. The fine-tuning objective is formulated as:
$\begin{split}\hat{\epsilon}^{*}_{\theta}(\mathbf{z}_{t},t,\mathbf{c})&=\epsilon_{\theta}(\mathbf{z}_{t},t)-w\cdot(\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{z}_{t},t))\\ &=\epsilon_{\theta}(\mathbf{z}_{t},t)+w\cdot(\epsilon_{\theta}^{*}(\mathbf{z}_{t},t,\mathbf{c})-\epsilon_{\theta}(\mathbf{z}_{t},t))\end{split}$ (4)
From this, we derive the unlearning loss like ESD [29]:
@@ III-C Key Step Unlearning Optimization
where $\epsilon^{*}_{\theta}$ represents the unlearned model, $\epsilon_{\theta}$ is the guidance model, $z_{t}$ denotes the denoised data at step $t$, and $c$ corresponds to the text prompt.
Unconditional Predicted Noise Optimization. More importantly, to the best of our knowledge, this is the first work to incorporate unconditional noise prediction into the optimization scope of concept unlearning. Initially, our goal was to constrain changes in the model’s unconditional prediction during training, aiming to preserve its generative performance. However, we identified a critical issue: although the prior of unconditional noise prediction is only conditioned on $\mathbf{z}_{t}$ and $t$, when $\mathbf{z}_{t}$ contains sufficient semantic information about the target concept $c$, the model’s unconditional prediction inevitably drifts toward $c$. To verify this, we empirically perform sampling using the prompt “naked woman”. At different denoising steps, we switch to purely unconditional noise prediction for the remaining steps, and apply NudeNet [59] to detect and score nudity in each generated image. As shown in Figure 4, even switching to unconditional sampling after just 5 standard steps produces images containing explicit nudity content. Furthermore, when switching after 30 steps, the generated results are almost indistinguishable from those produced by the full 50-step sampling. This indicates that, as the reverse process progresses, the sampling direction of unconditional noise prediction gradually shifts toward the concept $c$.
Therefore, we argue that unconditional prediction should also be partially optimized for concept unlearning. We propose the following regularization loss:
$\text{Loss}_{\text{regular}}=\left\|\epsilon_{\theta}^{*}(\mathbf{z}_{t},t)-\left((1+\lambda_{1})\epsilon_{\theta}(\mathbf{z}_{t},t)-\lambda_{1}\,\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})\right)\right\|_{2}^{2}$ (6)
Here, $\lambda_{1}$ is designed as a step-dependent coefficient to regulate the extent of concept unlearning applied to unconditional noise prediction. Since the bias toward concept $c$ increases with the number of reverse steps, we define $\lambda_{1}$ as a linear function with respect to $t$. Specifically, for $t$ ranging from 1000 to 0, we set $\lambda_{1}=(1000-t)*(2\times 10^{-5})$. The value was chosen empirically to balance semantic retention in early steps and suppression in later ones.
The final Key Step Unlearning Optimization function for KSCU is defined as:
where $\lambda_{2}=1\times 10^{-4}$ in KSCU.
III-D Training Algorithms
Algorithm 2 KSCU-Quick
Input: Unlearn Model $\epsilon^{*}_{\theta}$, Guidance Model $\epsilon_{\theta}$, scheduler $S$, target concept $c$, augmentation steps $N$, training steps $M$, Key Step Table
Output: $\epsilon^{*}_{\theta}$
1: Sample noise $n\sim\mathcal{N}(0,1)$
[…]
4:for $i=2$ to $M$ do
5:  if $i>N$ then
6:   $c^{*}\leftarrow\text{augment}(c)$
7:  else
8:   $c^{*}\leftarrow c$
9:  end if
[…]
11:  if $i<N$ then
12:   $t_{next}$ = $KeyStepTable[i+1]$
13:   if $t_{next}>t$ then
14:    $z_{t_{next}}=D(\epsilon^{*}_{\theta},z_{t},t_{next},t,c)$
15:   else
16:    $z_{t_{next}}=D(\epsilon^{*}_{\theta},n,t_{next},0,c)$
17:   end if
18:   $z_{t}=z_{t_{next}}$
19:  end if
20:end for
Although our KSCU framework significantly reduces the number of training iterations compared with conventional concept erasure approaches, the improvement in wall-clock training time remains limited. This is mainly because KSCU requires generating unlearning samples on-the-fly during fine-tuning, rather than pre-constructing the dataset as in other methods. Nevertheless, we argue that this online data generation strategy is crucial for the effectiveness of KSCU and thus remains an integral part of our design. On the other hand, we further observed that, benefiting from the step selection strategy in Algorithm 1, the steps chosen in the KeyStepTable exhibit a periodic and incrementally increasing pattern. Inspired by this observation, we propose an acceleration strategy that leverages such regularity to skip redundant intermediate sampling steps and only perform noise propagation and state updates at critical positions. This design substantially speeds up the training process while maintaining comparable performance to the original method. Concretely, the acceleration strategy reuses forward trajectories and recomputes states only at periodic checkpoints, effectively eliminating the computational overhead of repeated step-wise sampling. The complete KSCU algorithm, including this acceleration strategy, is summarized in Algorithm 2.
IV Experimental Results
TABLE II: Class and Style Unlearning Quantitative Performance on Unlearn Canvas.
class style
[…]
ESD [29] 0.81 68.6 98.9 96.4 26.1 0.91 100.0 85.7 99.0 25.3
RECE [26] 0.70 62.8 88.9 96.3 25.9 0.00 12.0 97.6 98.4 20.9
DUO [60] 0.00 27.8 89.9 75.2 95.8 0.00 31.5 77.0 89.8 69.1
CA [22] 0.08 32.5 86.7 83.4 51.9 0.63 96.0 86.6 88.7 37.4
FMN [24] 0.00 25.9 93.0 96.1 36.8 0.28 41.5 95.1 91.3 28.5
ANT [61] 0.21 44.1 80.7 89.6 54.7 0.01 14.5 91.7 78.2 45.6
EDiff [27] 0.68 78.0 94.1 90.1 48.2 0.38 67.5 93.7 98.3 39.8
Salun [23] 0.33 51.3 93.5 94.6 47.9 0.04 17.5 96.3 95.2 34.8
SHS [48] 0.16 62.5 54.6 56.7 79.4 0.04 65.0 51.1 29.4 65.9
KSCU 0.84 70.0 99.0 97.1 25.8 0.95 100.0 88.7 99.3 23.1
IV-A Experimental Setup
@@
Dataset. We evaluate concept unlearning across four task types—Class, Style, Instance, and Nudity—on two datasets: I2P [62] and Unlearn Canvas [63].
I2P contains 4,703 NSFW prompts. We selected 931 prompts labeled as “sexual” to generate images, from which Nudenet [59] (threshold = 0.2) detected 575 nudity regions. Across the full dataset, 793 nudity regions were found. This subset retains most of the evaluation potential while reducing computational cost.
Unlearn Canvas is a high-resolution stylized image dataset with 20 classes and 60 styles. Due to the high cost of full evaluation, we selected 10 diverse classes and 10 representative styles:
@@
  • Classes: Architectures, Bears, Cats, Dogs, Humans, Towers, Flowers, Sandwiches, Trees, Sea.
@@
Such selections also ensured diversity: the chosen classes include animals, plants, natural scenes, and man-made objects, while the styles cover both artistic styles, such as “Monet,” and illustrative styles, such as “Sketch”.
Evaluation. We adopt Unlearn Accuracy (UA) [29] as the evaluation metric, computed as $\text{UA}=1-x/y$. In I2P, $x$ is the number of nudity regions detected post-unlearning, and $y=575$. Unlike prior methods [29] that report the number of nude parts, we aggregate all detections into a single UA score for clarity and comparability. In Unlearn Canvas, $x$ is the number of classifier-detected target concepts, and $y$ is the number of generated images for the corresponding prompts.
@@
We also report Fréchet Inception Distance (FID) [64] to measure image quality. For I2P, prompts are drawn from COCO [65], and FID is computed against real images from coco30k. For Unlearn Canvas, images are generated with the prompt “A {class} image in {theme} style,” and compared to real images in the dataset. For Unlearn Canvas, we further report In-domain Retain Accuracy (IRA) and Cross-domain Retain Accuracy (CRA) to assess the preservation of non-target information. To evaluate defense performance, we generate adversarial prompts using UDA [54], reporting Attack Success Rate (ASR) on I2P.
More importantly, to intuitively compare the overall performance of different methods in terms of unlearning effectiveness and generative retainability, we introduce a statistical metric, $\text{Overall}=\text{UA}_{0\text{-}1}\times(1-\text{FID}_{0\text{-}1}),$ where $\text{UA}_{0\text{-}1}$ and $\text{FID}_{0\text{-}1}$ denote the min-max normalized UA and FID. Only methods that achieve both effective unlearning and high-generative retainability yield high overall scores, making it a direct indicator of performance trade-off.
Figure 5: KSCU’s Qualitative Performance on Unlearn Canvas. Red-bordered images show the generations of the unlearned concept after KSCU, while borderless images depict the results of unrelated concepts. More qualitative results and controlled replacement experiments are provided in Appendix A.
Baselines and Training Details. We compare KSCU with nine SOTA methods: ESD [29], RECE [26], CA [22], FMN [24], ANT [61] DUO [60], EDiff [27], Salun [23], and SHS [48]. All baselines use default hyperparameters from their original papers.
@@
KSCU was trained with a batch size of 1 and 50 timesteps. For style unlearning, the Key Step Table started at $S=25$ with $loop_{n}=8$, updating only the cross-attention modules for 500 optimizer iterations. For other tasks, it started at $S=15$ with $loop_{n}=8$, updating all modules except cross-attention in the U-Net for 700 iterations. We used the Adam optimizer with a learning rate of $1\times 10^{-5}$. All experiments employed DDIM with 50 denoising steps. Stable Diffusion v1.5 was used for Unlearn Canvas, while all other experiments (unless otherwise specified) used v1.4. Results for more versions of SD are included in the Appendix. All experiments were conducted on a single NVIDIA L40 GPU.
IV-B Class and Style Unlearning
@@
To evaluate the effectiveness of our proposed KSCU, we first performed qualitative experiments on class and style unlearning tasks. As shown in Figure 5, for different image samples, KSCU has completely blocked the generation of the target class or style (red-bordered images), with minimal impact on unrelated classes and styles (borderless ones). We then quantitatively evaluate its performance as follows.
Class Unlearning. As shown in Table II, KSCU achieves the highest overall score, indicating superior class unlearning while better preserving generative capability—an advantage not observed in other SOTA methods.
@@
Specifically, as illustrated in Table II, KSCU achieved 70%, 99.0%, 97.1%, and 25.8 on UA, IRA, CRA, and FID, respectively. In terms of UA, KSCU ranked as the second-best method, while it achieved the best performance on IRA, CRA, and FID. A method that aggressively removed target concepts might achieve higher UA but often at the cost of severely compromising image quality and diversity. Conversely, an overly conservative method may ensure generative retainability but fail to achieve effective unlearning. Therefore, evaluating a T2I diffusion model unlearning method requires consideration of both unlearning effectiveness (UA) and generative retainability (IRA, CRA, FID). In these two regards, KSCU demonstrates a more favorable trade-off. A high UA of 70% indicates strong unlearning effectiveness, while the best IRA and CRA scores confirm its superior ability to preserve generative diversity. Furthermore, KSCU outperforms all other SOTAs in FID, indicating that its unlearned model can generate higher-quality images, further validating its generative retainability. Conversely, while EDiff attained the highest UA, its FID reached 48, suggesting that although it exhibited strong unlearning performance, it significantly degraded the model’s generative ability. The low generative retainability of EDiff makes it difficult to use in practical applications.
@@
Style Unlearning. As shown in Table II, KSCU demonstrates strong unlearning effectiveness and retainability on the style unlearning task, achieving the highest overall score of 0.95. Specifically, KSCU achieves scores of 100%, 88.7%, 99.3%, and 23.1 in UA, IRA, CRA, and FID respectively, as shown in Table II. Notably, KSCU achieves the best performance in UA and CRA among all compared methods. While its FID and IRA are slightly lower than those of RECE, the latter exhibits extremely poor unlearning efficiency (UA == 12.0%), achieving higher IRA or FID at the cost of insufficient concept forgetting. In contrast, compared to high-efficiency unlearning methods like ESD (UA >> 95%), KSCU achieves significantly better results in IRA, CRA, and FID. These results suggest that KSCU not only effectively removes stylistic traces but also better preserves the model’s ability to generate unrelated concepts.
@@
By selectively fine-tuning only the last 50% of the denoising steps, KSCU reduces the training iterations by half compared to ESD. Despite the substantial reduction in training cost, KSCU still outperforms ESD in style unlearning, further demonstrating that full-step fine-tuning is unnecessary for concept unlearning in T2I diffusion models. Focusing on the most influential steps enables more efficient and controllable unlearning while minimizing disruption to the model’s generative capabilities. These findings indicate that the key step table optimization strategy adopted by KSCU offers a more practical and scalable solution for concept unlearning in large-scale diffusion models.
@@
TABLE III: Evaluation of Unlearning Capability on “Nudity”. The bold font represents the best, and the underlined font represents the second-best. Time is the first run time. Some methods include time to generate the required data.
[…]
ESD 0.71 88.2 73.7 15.4 2557
RECE 0.66 85.4 87.5 15.0 22
Salun 0.00 100.0 4.2 67.9 581
DUO 0.00 56.2 98.3 15.2 628
ANT 0.78 99.9 5.3 25.4 3016
EDiff 0.77 93.9 52.5 19.8 1914
SHS 0.59 84.3 95.5 18.1 817
KSCU 0.92 96.5 47.2 14.1 546
IV-C Nudity Unlearning & Robustness
Figure 6: Comparison of visualization results of forgetting four groups of instances using various concept unlearning methods. It can be observed that our method(KSCU) not only achieves more thorough instance forgetting but also better preserves unrelated generative quality.
Nudity Unlearning. In the NSFW concept unlearning experiment, we fine-tuned models to forget the concept of “nudity” and generated unlearned models. Each method was evaluated by generating 931 and 30,000 images using prompts from the i2p [62] and COCO [65] datasets, respectively. We compare unlearning effectiveness using Unlearn Accuracy and ASR, and assess generative retainability with FID. We also report the training time of each method to evaluate efficiency.